Prev Next

SW&IT Meeting Minutes 2022-04-04

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20220321
LR20220321 - Bec, Matthieu 80.0
LR20220321 - Mechtley, Brandon 80.0
LR20220321 - Eychaner, Glenn 76.0
LR20220321 - Edmeades, David 80.0
LR20220321 - Grenz, Paul 20.0
LR20220321 - Hooper, Stephen 0.0
LR20220321 - Zhang, Yang 80.0
LR20220321 - Zhang, Xianyu 80.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • SHARK TCS-sim/ALH answer meetup
    • IT8353, anon ftp -> https with account
    • refmemrecorder with JHill
  • This week
    • ATT, AO tests planning meeting (Mon @10am)
    • ARGOS github synch (Lorenzo?)
    • aorev_git dx asm, others ctd.
    • dms zwgd refactoring
    • refmemrecorder with JHill ctd?
    • TMS gmt/lbt round-up Fri.

YZ

  • Previous:
    • Worked on OSS/Argos proxy testing, but wasn't able to complete it (due to lack of telescope time).
    • Working on talk at the TMS workshop.
    • Met with Olga and discussed several issues (also initiated discussion with OSU team on one of the MODS issues).
    • Met with SHARK-NIR people on Friday, helping them with their dev efforts.
    • Fixed (hopefully) the recent TCS/IIF trouble with Argos swingarm.
    • Stephen helped fixing LBC CMU's NFS mount problem.
  • Next:
    • TMS workshop on Friday (still working on the talk).
    • Looking at some of the recent issues (mostly TCS).
    • Probably meet with Olga again on some LBC/MODS issues.

DE

  • Last week
    • Veeam licensing quote sorted, sent to purchasing. Should be done this week
    • Backup hiccups seem to have been resolved, no errors last weekend
    • LBTI storage--should be good to support them. The price on the expansion unit continues to fall in line with expectations.
    • Laptop shuffle
      • SH noted that the ex-ME laptop wouldn't power on; I will check it out next time I'm in.
    • No news re: either JW or AC new laptops.
    • Re-test Vault backup method, tests proved successful so we can ditch the ADMS database dump and parallel backups.
  • This week
    • Install new Veeam licenses when they arrive
    • Gently query about new laptops to make sure all is well
    • Got a lovely 4-week notice from Silva re: VPN user requests, will take care of that this week and have all of them test their accounts
    • Refresh my workstation's battery

GE

  • Last week:
    • Rewrote the document regarding migration of the LUCI observer interface from Java/Swing to HTML/CSS/Javascript.
    • Read TCS documentation on GitHub and in the Wiki, and started poking around my TCS VM to upgrade from 2018A to the latest version.
    • Further investigation of changes in RMI (Activation) and XML (JAXB) support in JVMs after version 8.
    • Started working on ITs related to the Observer Scheduler GUI (8571, 8529).
  • This week:
    • Upgrade my TCS VM to the latest version.
    • Work on Observer Scheduler GUI (8571, 8529).

BM

  • Last week:
    • set_z0 tests with Doug: nothing timed out @ 1200Hz, both banks gain=0 and gain!=0
    • bit of help w/Fabio setting up GH remote (couldn't get his deploy key working, so he just used https)
    • gzip log files on dxsoul--should be ok for a while
    • new issue: reset optical gain to 1 when autogain has loopfault--got feedback from Alfio, need to clarify with Juan Carlos/Jenny
    • possible new issue: AbstractSystem:_cmdResult thread safety (may be crashing when several commands sent rapidly)
    • tracked down NCPA_TRACK RTDB variable. Just need to add it to AOS.
    • some more work with Juan Carlos on interrupted offsets--appears disturbances stay on briefly after AdSec pause before Wfs pause--may need to swap order during AO pause. Not sure if this is related to the interrupted offsets, though. * This week:
  • This week:
    • to test: gain ramp
    • investigations:
      • soul-sxwfs BCU timeouts
      • interrupted offsets
      • msgd disconnect
    • ncpa status rtdb (?)
    • start looking at zeromq, dms

SH

  • IT meeting
  • LBTI questions
  • LBC NFS kernel module weirdness
  • Wiki problems
  • Mountain WiFi controller problems
  • Macbook for Pat
  • Sick Friday

next:

  • LBTI data
  • Core switches
  • VMware
Topic revision: r11 - 04 Apr 2022, BrandonMechtley
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback